THE OPTICAL SPECTRUM OF ATOMIC CLUSTERS Si_{3} and S_{3}

\underline{DAMIAN L. KOKKIN}, NEIL J. REILLY, MICHAEL C. McCARTHY, and PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138; JOHN F. STANTON, Department of Chemistry & Biochemistry, The University of Texas at Austin, 1 University Station A5300 Austin, TX 78712.

The optical spectra of the atomic clusters Si_{3} and S_{3} will be discussed in detail. The jet-cooled 2-colour resonant-2-photon ionisation spectrum of Si_{3} has now been detected for the first time in the 530 nm region. For Si_{3} a progression in the excited state bending and stretching modes are seen built onto the origin. Lower state frequencies derived from sequence band structure agree with those previously measured for the D_{3h} low lying electronic state. For thiozone additional experiments were undertaken and progressions in the excited state bending and stretching modes are seen with frequencies of 350 cm^{−1} and 420 cm^{−1} built onto the origin band at 23053.5 cm^{−1} respectively. In this talk our results are compared to theoretical results on both the C_{2v} and D_{3h} forms.